Metallica Bearings, Inc., is a young start-up company. No dividends will be paid on the stock over the next eleven years, because the firm needs to plow back its earnings to fuel growth. The company will then pay a dividend of $12.50 per share 12 years from today and will increase the dividend by 5.5 percent per year thereafter. If the required return on this stock is 13.5 percent, what is the current share price?
Answer #1

Year 12 dividend = 12.5

Year 13 dividend = 12.5 * 1.055 = 13.1875

Value at year 12 = D13 / require rate - growth rate

Value at year 12 = 13.1875 / 0.135 - 0.055

Value at year 12 = 13.1875 / 0.08

Value at year 12 = 164.84375

Current share price = 12.5 / (1 + 0.135)12 + 164.84375 / (1 + 0.135)12

Current share price = 2.735015 + 36.068

Current share price = $38.80